Two point charges reacted with each other to produce 4.32 x 10⁻⁴ J. the magnitude of q


is 27 nC with a distance from the point charge of 19 cm. Find the magnitude of the point


charge

1
Expert's answer
2022-03-16T18:28:33-0400

The potential energy of the system of charges is given by

"E_p=k\frac{q_1q_2}{r}"

Hence

"q_2=\frac{E_pr}{kq_1}"

"q_2=\frac{4.32*10^{-4}*0.19}{9*10^9*27*10^{-9}}=3.4*10^{-7}\:\rm C"